OpenStudy (anonymous):

Megan’s backyard has a small pond that is 15 feet by 10 feet. Her mother wants her to add a wooden deck pathway around the pond, so that the total area of the pool and deck is 234 ft2. How wide does the pathway need to be to fulfill Megan’s mom’s request?

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

Pond and pathway: Length = 15 + x + x = 15 + 2x = 2x + 15 Width: = 10 + x + x = 10 + 2x = 2x + 10 Area = LW = (2x + 15)(2x + 10) Total area = 234 Solve (2x + 15)(2x + 10) = 234 FOIL left side. Subtract 234 to left side. Collect like terms. Try to factor left side. If not, then use the quadratic formula.